If a^2+b^2+c^2=250 and ab+bc not equal to ca, find a+b+c

Grade:9

1 Answers

Arun
5 years ago
 
(a + b + c)^2 = a^2 + b^2 + c^2 + 2 (ab + bc + ca)
 
a + b + c = 250 + 2 (ab + bc + ca)
 
further I didnot understand the condition given by you.
Please chck the question
 
Regards
Arun